Multiply, Simplify Completely
1/84 + 1/60

Found 2 solutions by Narsinghchauhan8858795971, solver91311:
Answer by Narsinghchauhan8858795971(1)

1/84 + 1/60
solution: LCM of 84 and 60 is 420
so 1/84 + 1/60 = 5+7/420
= 12/420
= 1/35

Answer by solver91311(24713)



Multiply? Why do you want to multiply when this is an addition problem?

John

My calculator said it, I believe it, that settles it